# On additive error approximations to ##\##BQP

### v1

@@ Abstract
Counting complexity characterizes the difficulty of computing functions related to the number of valid certificates to efficiently verifiable decision problems. Here we study additive approximations to a quantum generalization of counting classes known as #BQP. First, we show that there exist efficient quantum algorithms that achieve additive approximations to #BQP problems to an error exponential in the number of witness qubits in the corresponding verifier circuit, and demonstrate that the level of approximation attained is, in a sense, optimal. We next give evidence that such approximations can not be efficiently achieved classically by showing that the ability to return such approximations is BQP-hard. We next look at the relationship between such additive approximations to #BQP and the complexity class DQC1, showing that a restricted class of #BQP problems are DQC1-complete.
@@ 1 Introduction
The class #P precisely characterizes the complexity of counting problems such as computing the permanent of binary matrices, counting the number of perfect matchings in a graph [26], and computing partition functions of certain classical spin systems [17, 6]. On the other hand, a number of important problems in, e.g., algebraic combinatorics [7] and quantum computing [5] are polynomial-time reducible to #P but are not #P problems. The closure of #P under subtraction defines the class GapP [12] which is known to contain problems such as computing multiplicities arising in representation theory of the symmetric group [7].
@@
The counting class #BQP was defined to formalize the notion of counting the number of quantum witnesses to promise problems that allow efficient verification on quantum computers [5, 21]. This class contains quantum-inspired counting problems such as computing the density of states of quantum Hamiltonians [5] and computing representation theoretic multiplicities such as Kronecker coefficients [16, 3]. Though #P, GapP, and #BQP are equivalent under polynomial-time reductions [5, 11], the subtle difference amongst them manifests directly in the hardness of approximation, specifically approximating the respective quantities to within a multiplicative error. Any #P function can be multiplicatively approximated in the third level of the polynomial hierarchy [24], a class of problems believed to be easier than #${\sf\#P}$, but approximating GapP functions can be #P-hard in general [14]. However, it is unknown if multiplicatively approximating #BQP relations is any easier even though this has been used to define a new class QXC of quantum approximate counting problems [3, 4].
@@
It is thus interesting to ask if other notions of approximation also lead to problems of varying complexity. The study of additive approximations to counting problems was initiated in Ref. [1], motivated by the result that certain additive approximations to GapP problems suffice to solve BQP decision problems [12]. There, it was shown that all #P and GapP functions have efficient classical additive approximation algorithms up to a normalization factor exponential in the size of the descriptions of the solutions, with no difference in the complexity of obtaining the additive approximations for #P and GapP. In quantum computation, additive error approximations to some #BQP and QXC problems have been studied in the context of DQC1, the class of decision problems that can be solved in a model of quantum computation where the system is initialized with only one pure qubit [18, 20]. Specifically, decision problems based on certain additive error approximations to the density of states and quantum partition functions have been shown to be complete for DQC1 [2, 9, 8].
@@
In this work, we study the complexity of obtaining additive error approximations to #BQP problems and its connection to the complexity classes BQP and DQC1. We prove that there exists an efficient quantum algorithm which produces additive approximations to all #BQP problems up to a normalization exponential in the the number of qubits of the quantum witness. We then give evidence that the level of approximation achieved quantumly is not achievable by any polynomial-time classical algorithm by showing that it is BQP-hard. Next, we show that this level of approximation is likely optimal as the problem of computing more accurate approximations becomes #BQP-hard. We demonstrate, however, that a weaker additive approximation is achievable via a polynomial-time classical algorithm. Next we formalize results that elucidate the connection between DQC1 and #BQP. We show that DQC1 is equivalent to computing additive approximations to a subset of #BQP problems, namely those with verifier circuits containing at most a logarithmic number of ancilla qubits.
2 Background and Summary of Results
@@ 2.1 Classical Counting Complexity
Classical counting complexity, including classes such as #P and GapP, aims to characterize the difficulty of computing functions related to the number of accepting witnesses to NP problems. A decision problem is in NP if given an input $x$ that is a yes instance, there exists an efficiently generated polynomial-time verification procedure and at least one witness $y$ such that the procedure will accept on the input $(x,y)$, while for no instances, no such witness exists. The class #P is the set of functions that can be defined by counting the number of witnesses to an NP problem.
We can conveniently define both NP and #P via relations $R\subset\{0,1\}^{*}\times\{0,1\}^{*}$ between the set of inputs and the set of possible solutions such that there exists a Turing machine that decides the proposition $R(x,y)$ in polynomial-time. Introducing the set of accepting witnesses (of length bounded by a polynomial $p$ of the input) for input $x$,
@@
solving the NP decision problem for input $x$ reduces to deciding whether $|R(x)|>0$ while solving the #P problem for input $x$ is to return $|R(x)|$. The class GapP is then defined as the closure of #P under subtraction. Namely, a function is in GapP if and only if it can be expressed as the difference of two #P functions.
2.2 Acceptance Operators, QMA, and DQC1
A quantum generalization of NP is the class QMA, which roughly corresponds to NP except the witnesses are quantum states and the verification procedure is a quantum circuit. We will define QMA, and later the quantum analogue of #P, known as #BQP, via properties of acceptance operators $\mathcal{V}$ [4, 5] which generalize the relations $R$ in the previous section. We define them via a quantum circuit $V$ that acts on a tensor product Hilbert space $\mathcal{H}_{I}\otimes\mathcal{H}_{A}\otimes\mathcal{H}_{W}$ where $\mathcal{H}_{W}$ is the subsystem corresponding to the quantum witnesses, $\mathcal{H}_{A}$ corresponds to the subsystem of the ancillas needed for the verification procedure and $\mathcal{H}_{I}$ is the subsystem corresponding to the (classical) input. Given the circuit $V$ requiring $a$ ancilla qubits and a classical input bit string $x$, we define the acceptance operator $\mathcal{V}_{x}$ via
$\mathcal{V}_{x}=(\bra{x}\otimes\bra{0^{a}}\otimes\mathbbm{I}_{W})V^{\dagger}\Pi_{\text{out}}V(\ket{x}\otimes\ket{0^{a}}\otimes\mathbbm{I}_{W})\,,$ (2)
where $\Pi_{\text{out}}$ is a projector onto the |$\outerproduct{1}{1}$ | state of the first qubit. With this we define the complexity class QMA.
@@ 2.2 Acceptance Operators, QMA, and DQC1
    if $x\in A_{\text{YES}}$ then $||\mathcal{V}^{(|x|)}_{x}||\geq c(|x|)$ ,
@@
    if $x\in A_{\text{NO}}$ then $||\mathcal{V}^{(|x|)}_{x}||\leq s(|x|)$ ,
where $\mathcal{V}^{(|x|)}_{x}$ is the acceptance operator of Eq. (2) corresponding to $V^{(|x|)}$ on input $x$.
Often in the literature the functions $c$ and $s$ are taken to be the constants $2/3$ and $1/3$, respectively. We choose this definition as it motivates the discussion of the promise gap $c(n)-s(n)$, which becomes important when defining #BQP in Section 2.3.
We will now recall the standard definition of BQP, and then afterwards define a complete problem for BQP that can be expressed in terms of properties of the acceptance operators.
Definition 2 (BQP).
  1. 1.
    if $x\in A_{\text{YES}}$ then $\bra{x,0^{a}}V^{(|x|)\dagger}\Pi_{\textup{out}}V^{(|x|)}\ket{x,0^{a}}\geq c(|x|)$ ,
  2. 2.
    if $x\in A_{\text{NO}}$ then $\bra{x,0^{a}}V^{(|x|)\dagger}\Pi_{\textup{out}}V^{(|x|)}\ket{x,0^{a}}\leq s(|x|)$
The BQP-complete problem we introduce related to verifier circuits is the decision problem of estimating their average accept probability.
Definition 3 (Average Accept Probability of Verifier Circuits).
  1. 1.
    $\frac{\text{Tr}[\mathcal{V}^{(|x|)}_{x}]}{2^{w(|x|)}}\geq c(|x|)$ ,
  2. 2.
    or $\frac{\text{Tr}[\mathcal{V}^{(|x|)}_{x}]}{2^{w(|x|)}}\leq s(|x|)$ ,
promised one of these is the case, where $\mathcal{V}^{(|x|)}_{x}$ is the acceptance operator of Eq. (2) corresponding to $V^{(|x|)}$ on input $x$.
We show in Appendix B that this problem is BQP-complete.
[…]
$\mathcal{D}_{x}=(\bra{0^{a}}\otimes\mathbb{I}_{W})D_{x}^{\dagger}\Pi_{\text{out}}D_{x}(\ket{0^{a}}\otimes\mathbb{I}_{W})\,.$ (3)
With this we formally define DQC1.
Definition 4 (DQC1).
Let $A=(A_{\text{YES}},A_{\text{NO}})$ be a promise problem and let c,s: $\mathbb{N}\to[0,1]$ be functions satisfying $c(n)-s(n)=\Omega(\text{poly}(n)^{-1})$. Then $A\in{\sf DQC}_{1}$ if and only if for every input $x\in A$ there exists a polynomial-time generated quantum circuit $D_{x}$ with corresponding acceptance operator $\mathcal{D}_{x}$ as in Eq. (3) acting on $w(|x|)=O(\text{poly}(|x|))$ witness qubits and $a(|x|)=O(log(|x|))$ ancilla qubits such that:
  1. 1.
    If $x\in A_{\text{YES}}$ then $\frac{\textup{Tr}[\mathcal{D}_{x}]}{2^{w(|x|)}}\geq c(|x|)$ ,
  2. 2.
    If $x\in A_{\text{No}}$ then $\frac{\textup{Tr}[\mathcal{D}_{x}]}{2^{w(|x|)}}\leq s(|x|)$
In the literature DQC1 is often alternatively viewed as the class of problems that can be solved via sampling from circuits of the form given above. A canonical complete problem from this point of view is that of estimating the trace of $2^{n}\times 2^{n}$ unitary matrices up to an additive error normalized by $2^{n}$ (see Definition 6). Problems that are not decision problems are often referred to as DQC1-complete if they are polynomial-time reducible to this problem. We will adopt a similar point of view in this paper, although when showing DQC1-hardness of a problem we will only show that it holds for the decision problem version, noting that the use of a binary search can establish polynomial-time reducibility between the decision problem version of DQC1 and the problem of additively estimating traces of unitary matrices to the specified normalization.
@@
Some care needs to be taken in defining the counting class associated to QMA, known as #BQP. Intuitively, one would like to generalize the notion of computing the size of the set $R(x)$ (1) to the quantum case. This is not straightforward, as the number of possible quantum witnesses is uncountably infinite. Instead, a counting problem can be formulated as determining the dimension of the subspace of good witnesses. This subspace will be defined via the spectral properties of the acceptance operator $\mathcal{V}_{x}$. We first define the maps $N_{\geq a}$ from Hermitian operators to integers via
@@
That is, they return the number of eigenvalues of $H$ that are greater than or equal to $a$. One can then attempt to define #BQP via counting functions such that given a circuit $V$ and an input $x$ corresponding to an acceptance operator $\mathcal{V}_{x}$, return $N_{\geq c}(\mathcal{V}_{x})$. This ends up being too powerful, as the ability to return this would allow one to solve problems in the variant of QMA where the promise gap, $c-s$, is allowed to be exponentially small, which characterizes the class PSPACE [10]. Thus, one must be careful with how the promise gap is handled.
The original formulation of #${\sf\#BQP}$ [5, 21] circumvents this by imposing a promise that acceptance operators $\mathcal{V}_{x}$ corresponding to allowed problem instances $x$ do not have any eigenvalues in the interval $(c,s)$. Conveniently, this also guarantees that the output of any #${\sf\#BQP}$ function is unique but has the consequence that proving inclusion of problems in #${\sf\#BQP}$ requires adding extra promises. In particular, Ref. [5] showed that the density of states is a #${\sf\#BQP}$-complete problem but only for Hamiltonians that satisfy a potentially nonphysical promise on their spectrum. As Ref. [5] notes, there is an alternative approach to defining #${\sf\#BQP}$ where the output is allowed to “miscount” eigenvalues in the interval $(c,s)$. This was later generalized to define the notion of relative-error approximations to #${\sf\#BQP}$ problems and to show that this is equivalent to problems such as computing relative-error approximations to quantum partition functions under polynomial-time reductions [4]. Here, we choose to work with the miscounting definition of #${\sf\#BQP}$ because it does not require extra assumptions on the associated verifier circuits.
In order to rigorously define #BQP problems with the allowed miscounting in the promise gap we will define them as relations $r\subset J\times\mathbb{Z}^{+}$ between a set of inputs $J\subset\{0,1\}^{*}$, and the set of positive integers $\mathbb{Z}^{+}$. We use the lower case $r$ to distinguish them from the relations $R$ we used to define #P. Note that this definition includes standard functions as a special case of relations satisfying the property that every $x\in J$ has exactly one $y\in\mathbb{Z}^{+}$ related to it. We denote the set of outputs related to the input $x\in J$ via
@@
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if there exists a uniformly generated family of quantum circuits $V^{r,n}$ and functions $c^{r},s^{r}:\mathbb{N}\to[0,1]$ satisfying $c^{r}(n)-s^{r}(n)=\Omega(\text{poly}(n)^{-1})$ such that
$r(x)=[N_{\geq c^{r}(|x|)}(\mathcal{V}_{x}^{r,|x|})..N_{\geq s^{r}(|x|)}(\mathcal{V}_{x}^{r,|x|})]\,.$
For the purposes of this work, certain parameters of the circuits $V^{r}_{x}$ will be very important. These parameters are defined as follows: given a family of verifier circuits $\{V^{r,n}\}_{n\in\mathbb{N}}$ in the $\{\text{Toffoli},H,S\}$ gate set where $r$ denotes the family and $n$ denotes the input size, define the parameters $w_{r},a_{r},t_{r},h_{r}:\mathbb{Z}^{+}\to\mathbb{Z}^{+}$ for a given input size $n$ via
@@
2.4 Additive Error Approximations
In Ref. [1] it was shown that all #P and GapP functions have a natural level of additive approximability related to the length of the witnesses of the associated NP problems defining them. In particular, given a #P or GapP function $f$ with witnesses on $w(|x|)$ bits for inputs of size $|x|$, the authors show that there exists a probabilistic algorithm on input $x$ that outputs an estimate $\tilde{f}(x)$ satisfying

### v2

# Quantum Approximate Counting with Additive Error: Hardness and Optimality

@@ Abstract
Quantum counting is the task of determining the dimension of the subspace of states that are accepted by a quantum verifier circuit. It is the quantum analog of counting the number of valid solutions to NP problems - a problem well-studied in theoretical computer science with far-reaching implications in computational complexity. The complexity of solving the class #BQP of quantum counting problems, either exactly or within suitable approximations, is related to the hardness of computing many-body physics quantities arising in algebraic combinatorics. Here, we address the complexity of quantum approximate counting under additive error. First, we show that computing additive approximations to #BQP problems to within an error exponential in the number of witness qubits in the corresponding verifier circuit is as powerful as polynomial-time quantum computation. Next, we show that returning an estimate within error that is any smaller is #BQP-hard. Finally, we show that additive approximations to a restricted class of #BQP problems are equivalent in computational hardness to the class DQC1. Our work parallels results on additively approximating #P and GapP functions.
@@ 1 Introduction
The class #P precisely characterizes the complexity of counting problems such as computing the permanent of binary matrices, counting the number of perfect matchings in a graph [26], and computing partition functions of certain classical spin systems [17, 6]. On the other hand, a number of important problems in, e.g., algebraic combinatorics [7] and quantum computing [5] are polynomial-time reducible to #P but are not #P problems. Another important counting class is GapP [12], defined as the closure of #P under subtraction, which is known to contain problems such as computing multiplicities arising in the representation theory of the symmetric group [7].
@@
The counting class #BQP was defined to formalize the notion of counting the number of quantum witnesses to promise problems that allow efficient verification on quantum computers [5, 21]. This class contains quantum-inspired counting problems such as computing the density of states of quantum Hamiltonians [5] and computing representation theoretic multiplicities such as Kronecker coefficients [16, 3]. Though #P, GapP, and #BQP are equivalent under polynomial-time reductions [5, 11], the subtle difference between them manifests directly in the hardness of approximation, specifically in approximating the respective quantities within a multiplicative error. Any #P function can be multiplicatively approximated in the third level of the polynomial hierarchy [24], a class of problems believed to be easier than #${\sf\#P}$, but approximating GapP functions can be #P-hard in general [14]. However, it is unknown if multiplicatively approximating #BQP relations is any easier even though this has been used to define a new class QXC of quantum approximate counting problems [3, 4].
@@
It is thus interesting to ask if other notions of approximation also lead to problems of varying complexity. The study of additive approximations to counting problems was initiated in Ref. [1], motivated by the result that certain additive approximations to GapP problems suffice to solve BQP decision problems [12]. There, it was shown that all #P and GapP functions have efficient classical additive-approximation algorithms up to a normalization factor exponential in the size of the descriptions of the solutions, with no difference in the complexity of obtaining the additive approximations for #P and GapP. In quantum computation, additive-error approximations to some #BQP and QXC problems including estimating the quantum density of states and quantum partition functions have been shown to be complete for the class DQC1 [2, 9, 8], the set of decision problems that can be solved in a model of quantum computation where the system is initialized with only one pure qubit [18, 20].
@@
In this work, we study the complexity of obtaining additive-error approximations to #BQP problems and its connection to the complexity classes BQP and DQC1. We first prove that returning additive-error approximations to #BQP functions, normalized by the dimension of the witness subspace, is polynomial time equivalent to BQP, and thus captures the difficulty of polynomial-time quantum computation. Next, we show that this level of approximation is likely optimal as the problem of computing more accurate approximations becomes #BQP-hard. We then demonstrate that a weaker additive approximation is achievable via a polynomial-time classical algorithm. Next we present results that elucidate the connection between DQC1 and #BQP. We show that DQC1 is equivalent to computing additive approximations to a subset of #BQP problems—those with verifier circuits containing at most a logarithmic number of ancilla qubits.
2 Background
@@ 2.1 Classical Counting Complexity
Classical counting complexity, including classes such as #P and GapP, characterizes the difficulty of computing the number of accepting witnesses to NP problems. A problem is in NP if given a “yes” instance $x$ there is an efficiently generated polynomial-time verification procedure and at least one witness $y$ such that the procedure accepts on $(x,y)$, while for “no” instances, no such witness exists.
We choose to define both NP and #P via relations $R\subset\{0,1\}^{*}\times\{0,1\}^{*}$ between the set of inputs and the set of possible witnesses such that there exists a polynomial-time Turing machine that decides the proposition $R(x,y)$. Introducing the set of accepting witnesses of length bounded by a polynomial $p$ for input $x$:
@@
solving the NP decision problem is to decide whether $|R(x)|>0$ while solving the #P problem is to return $|R(x)|$. The class GapP is then defined as the set of functions that are the difference of two #P functions.
2.2 Acceptance Operators and QMA
A quantum generalization of NP is the class QMA, which roughly corresponds to NP except the witnesses are quantum states and the verification procedure is a quantum circuit. We will define QMA, and later the quantum analogue of #P, known as #BQP, via properties of acceptance operators $\mathcal{V}$ [4, 5] which generalize the relations $R$ in the previous section. We define them via a quantum circuit $V$ that acts on a Hilbert space $\mathcal{H}_{I}\otimes\mathcal{H}_{A}\otimes\mathcal{H}_{W}$ where $\mathcal{H}_{W}$ is the subspace spanned by the quantum witnesses, $\mathcal{H}_{A}$ is the ancilla subspace needed for the verification procedure, and $\mathcal{H}_{I}$ is the subspace of the (classical) input. Given the circuit $V$ requiring $a$ ancilla qubits and a classical input bit string $x$, we define the acceptance operator $\mathcal{V}_{x}$ via
$\mathcal{V}_{x}=(\bra{x}\otimes\langle 0^{a}|\otimes\mathbb{I}_{W})V^{\dagger}\Pi_{\text{out}}V(\ket{x}\otimes|0^{a}\rangle\otimes\mathbb{I}_{W})\,,$ (2)
where $\Pi_{\text{out}}$ is a projector onto the |$\outerproduct{1}{1}$ | state of the first qubit. Using these operators, we can define the complexity class QMA.
@@ 2.2 Acceptance Operators and QMA
    if $x\in A_{\text{YES}}$, then $||\mathcal{V}_{x}||\geq c(|x|)$ ,
@@
    if $x\in A_{\text{NO}}$, then $||\mathcal{V}_{x}||\leq s(|x|)$ ,
where $\mathcal{V}_{x}$ is the acceptance operator of Eq. (2) corresponding to $V^{(|x|)}$ on input $x$.
Often in the literature the functions $c$ and $s$ are taken to be the constants $2/3$ and $1/3$, respectively. We choose this definition as it motivates the discussion of the promise gap $c(n)-s(n)$, which becomes important when defining #BQP in Section 2.3 below.
@@
Some care needs to be taken in defining the counting class associated to QMA, known as #BQP. Intuitively, one would like to generalize the notion of computing the size of the set $R(x)$ in Eq. (1) to the quantum case. This is not straightforward, as the number of possible quantum witnesses is uncountably infinite. Instead, a counting problem can be formulated as determining the dimension of the subspace of good witnesses. This subspace will be defined via the spectral properties of the acceptance operator $\mathcal{V}_{x}$. We first define the maps $N_{\geq a}$ from Hermitian operators to integers via
@@
That is, they return the number of eigenvalues of $H$ that are greater than or equal to $a$. One can then attempt to define #BQP via counting functions such that given a circuit $V$ and an input $x$ corresponding to an acceptance operator $\mathcal{V}_{x}$, return $N_{\geq c}(\mathcal{V}_{x})$. This ends up being too powerful, as the ability to return this would allow one to solve problems in the variant of QMA where the promise gap, $c-s$, is allowed to be exponentially small. This instead characterizes the class PSPACE [10]. Thus, one must be careful with how the promise gap is handled.
The original formulation of #${\sf\#BQP}$ [5, 21] circumvents this by imposing a promise that acceptance operators $\mathcal{V}_{x}$ corresponding to allowed problem instances $x$ do not have any eigenvalues in the interval $(c,s)$. Conveniently, this also guarantees that the output of any #${\sf\#BQP}$ function is unique but has the consequence that proving inclusion of problems in #${\sf\#BQP}$ requires adding extra promises. In particular, Ref. [5] showed that the density of states is a #${\sf\#BQP}$-complete problem but only for Hamiltonians that satisfy a potentially nonphysical promise on their spectrum. As Ref. [5] notes, there is an alternative approach to defining #${\sf\#BQP}$ where one is allowed to “miscount” eigenvalues in the interval $(c,s)$. This was later generalized to define the notion of relative-error approximations to #${\sf\#BQP}$ problems and to show that this is equivalent to problems such as computing relative-error approximations to quantum partition functions under polynomial-time reductions [4]. Here, we choose to work with the definition of #${\sf\#BQP}$ which allows for miscounting, as it does not require additional assumptions on the associated verifier circuits.
In order to rigorously define #BQP problems with the allowed miscounting in the promise gap, we again define them as relations $r\subset J\times\mathbb{Z}^{+}$ between a set of inputs $J\subset\{0,1\}^{*}$, and the set of positive integers $\mathbb{Z}^{+}$. Note that the use of $r$ is to distinguish them from the relations $R$ used in the definition of #P. The relations considered here include standard functions as a specific instance, satisfying the property that for every $x\in J$ there is exactly one $y\in\mathbb{Z}^{+}$ related to it. We denote the set of outputs related to the input $x\in J$ via
@@
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if there exists a uniformly-generated family of quantum circuits $V^{r,(n)}$ and functions $c^{r},s^{r}:\mathbb{N}\to[0,1]$ satisfying $c^{r}(n)-s^{r}(n)=\Omega(\text{poly}(n)^{-1})$ such that
$r(x)=[N_{\geq c^{r}(|x|)}(\mathcal{V}_{x}^{r})..N_{\geq s^{r}(|x|)}(\mathcal{V}_{x}^{r})]\,,$
where $\mathcal{V}_{x}^{r}$ is the acceptance operator corresponding to the circuit $V^{r,(|x|)}$ on input $x$.
Here the notation $[a..b]$ denotes the set of integers between and including $a$ and $b$. For the purposes of this work, certain parameters of the circuits $V^{r,(n)}$ will be very important. These parameters are defined as follows: Given a family of verifier circuits $\{V^{r,(n)}\}_{n\in\mathbb{N}}$ in the $\{\text{Toffoli},H,S\}$ gate set where $r$ denotes the family and $n$ denotes the input size, define the parameters $w_{r},a_{r},t_{r},h_{r}:\mathbb{Z}^{+}\to\mathbb{Z}^{+}$ via
@@
2.4 BQP and DQC1
We now recall the standard definition of BQP, which is the class of decision problems that can be solved by polynomial-time quantum computation.
Definition 3 (BQP).
  1. 1.
    if $x\in A_{\text{YES}}$, then $\bra{x,0^{a}}V^{(|x|)\dagger}\Pi_{\textup{out}}V^{(|x|)}\ket{x,0^{a}}\geq c(|x|)$ ,
  2. 2.
    if $x\in A_{\text{NO}}$, then $\bra{x,0^{a}}V^{(|x|)\dagger}\Pi_{\textup{out}}V^{(|x|)}\ket{x,0^{a}}\leq s(|x|)$
[…]
$\mathcal{D}_{x}=(\langle 0^{a}|\otimes\mathbb{I}_{W})D^{(x)\dagger}\Pi_{\text{out}}D^{(x)}(|0^{a}\rangle\otimes\mathbb{I}_{W})\,.$ (5)
With this clarified, we are now in a position to define DQC1.
Definition 4 (DQC1).
Let $A=(A_{\text{YES}},A_{\text{NO}})$ be a promise problem and let c,s: $\mathbb{N}\to[0,1]$ be functions satisfying $c(n)-s(n)=\Omega(\text{poly}(n)^{-1})$. Then $A\in{\sf DQC}_{1}$ if and only if for every input $x\in A$ there exists a polynomial-time-generated quantum circuit $D^{(x)}$ with corresponding acceptance operator $\mathcal{D}_{x}$ as in Eq. (5) acting on $w(|x|)=O(\text{poly}(|x|))$ witness qubits and $a(|x|)=O(log(|x|))$ ancilla qubits such that:
  1. 1.
    If $x\in A_{\text{YES}}$, then $\frac{\textup{Tr}[\mathcal{D}_{x}]}{2^{w(|x|)}}\geq c(|x|)$ ,
  2. 2.
    If $x\in A_{\text{NO}}$, then $\frac{\textup{Tr}[\mathcal{D}_{x}]}{2^{w(|x|)}}\leq s(|x|)$
2.5 Additive-Error Approximations
In Ref. [1] it was shown that all #P and GapP functions can be additively approximated to accuracies related to the length of the witnesses of the associated NP problems defining them. In particular, given a #P or GapP function $f$ with witnesses on $w(|x|)$ bits for inputs of size $|x|$, the authors show that there exists a classical probabilistic algorithm on input $x$ that outputs an estimate $\tilde{f}(x)$ satisfying